# Asymptotically Stable Quaternion-valued Hopfield-structured Neural Networks with Periodic Projection-based Supervised Learning Rules

### v1

@@ Appendix F Baselines
I. Damped Least-Squares Inverse Kinematics (IK): A closed-loop control law that solves joint increments directly based on Jacobian matrices and can be run online in milliseconds. It ensures local asymptotic convergence, but it is strongly dependent on the initial point and often only achieves a low accuracy of rad.
 II. Covariant Hamiltonian Optimization for Motion Planning (CHOMP): Optimized trajectory planning to generate smooth paths, but requires offline parameter adjustment and is time-consuming online.
 III. Stochastic Trajectory Optimization for Motion Planning (STOMP): An iterative trajectory optimizer that samples noisy perturbations of an initial guess and weights them by smoothness and collision‑avoidance costs; while it generates low‑curvature paths, it depends heavily on the initial trajectory and requires tens to hundreds of milliseconds per planning episode, lacking global convergence guarantees.
 IV. RRT-Connect + B-Spline Smoothing: First, use a fast random tree (RRT-Connect) to quickly generate a feasible path, and then use B-Spline or quadratic programming to smooth the path. This not only retains the efficient connectivity of the sampling algorithm, but also obtains a certain degree of smoothing effect. However, the continuity and convergence lack strict guarantees.
@@ Appendix G Absolute error to the target
The rapid convergence of exact error between the target state and the network state is already guaranteed by exponential or asymptotic stability under the framework of Lyapunov theory. Here we supplement a deduction for a explicit expression. Recall the notations for the trajectory and target are $\boldsymbol{q}(t)\in\mathbb{C}^{2}([0,+\infty])^{n},\ \boldsymbol{q}_{d}\in\mathbb{R}^{n}$, and the square error is expressed by:

### v2

@@ Appendix F Baselines
I. Damped Least-Squares Inverse Kinematics (IK) [7]: A closed-loop control law that solves joint increments directly based on Jacobian matrices and can be run online in milliseconds. It ensures local asymptotic convergence, but it is strongly dependent on the initial point and often only achieves a low accuracy of rad.
 II. Covariant Hamiltonian Optimization for Motion Planning (CHOMP) [40]: Optimized trajectory planning to generate smooth paths, but requires offline parameter adjustment and is time-consuming online.
 III. Stochastic Trajectory Optimization for Motion Planning (STOMP) [14]: An iterative trajectory optimizer that samples noisy perturbations of an initial guess and weights them by smoothness and collision‑avoidance costs; while it generates low‑curvature paths, it depends heavily on the initial trajectory and requires tens to hundreds of milliseconds per planning episode, lacking global convergence guarantees.
 IV. RRT-Connect + B-Spline Smoothing [2]: First, use a fast random tree (RRT-Connect) to quickly generate a feasible path, and then use B-Spline or quadratic programming to smooth the path. This not only retains the efficient connectivity of the sampling algorithm, but also obtains a certain degree of smoothing effect. However, the continuity and convergence lack strict guarantees.
@@ Appendix G Absolute error to the target
The rapid convergence of exact error between the target state and the network state is already guaranteed by exponential or asymptotic stability under the framework of Lyapunov theory. Here we supplement a deduction for an explicit expression. Recall the notations for the trajectory and target are $\boldsymbol{q}(t)\in\mathbb{C}^{2}([0,+\infty])^{n},\ \boldsymbol{q}_{d}\in\mathbb{R}^{n}$, and the square error is expressed by: